Log of Important Operation Events and Happenings
2015
- 2015-05-26 14:48
DHU commands the Software Under Voltage Control (UVC) mode #5.
The mission instruments XIS and HXD were shutdown.
No scientific data were available until the recovery of the UVC.
- 2015-01-10 20:55
BCCU commands the Under Voltage Control (UVC) mode.
The attitude control system was automatically switched into the safehold mode.
The mission instruments XIS and HXD were also shutdown.
No scientific data were available until the recovery of the UVC.
- 2015-01-05
BAT-B: Trickle charging was restarted.
- 2015-01-03 07:59
BCCU commands the Under Voltage Control (UVC) mode.
The attitude control system was automatically switched into the safehold mode.
The mission instruments XIS and HXD were also shutdown.
No scientific data were available until the recovery of the UVC.
2014
- 2014-08-28
BAT-B: Charging was disabled
- 2014-09-04 00:11
BCCU commands the Under Voltage Control (UVC) mode.
The attitude control system was automatically switched into the safehold mode.
The mission instruments XIS and HXD were also shutdown.
No scientific data were available until the recovery of the UVC.
- 2014-08-28 04:27
BCCU commands the Under Voltage Control (UVC) mode.
The attitude control system was automatically switched into the safehold mode.
The mission instruments XIS and HXD were also shutdown.
No scientific data were available until the recovery of the UVC.
- 2014-08-20 -- 2014-08-25 (G306.3-0.9)
Failure of the star tracking by STT.
- 2014-05-15 20:33:44
BCCU commands the Under Voltage Control (UVC) mode.
The attitude control system was automatically switched into the safehold mode.
The mission instruments XIS and HXD were also shutdown.
No scientific data were available until the recovery of the UVC.
- 2014-4-29 23:12:00 ~ 2014-04-30 15:05:00 (809039010FJ0854+4908_OFFSET)
Failure of the star tracking by STT (EC 63).
- 2014-4-28 22:01:00 ~ 2014-04-29 23:12:00 (809038010FJ0854+4908)
Failure of the star tracking by STT (EC 63).
- 2014-03-26
ATOG for fake safehold is disabled.
- 2014-03-16 09:10:00 ~ 2014-03-21 02:15:00 (408029010V1159_ORI)
Failure of the star tracking by STT (EC 63).
- 2014-02-22 00:00:00
XRT heaters are shut-off during the earth shade.
- 2014-01-22
BAT-B: Trickle charging was started.
- 2014-01-13 ~ 2014-01-24
BCCU commands the Under Voltage Control (UVC) mode frequently.
- 2014-01-09 18:39:19
BCCU commands the Under Voltage Control (UVC) mode.
The attitude control system was automatically switched into the safehold mode.
The mission instruments XIS and HXD were also shutdown.
No scientific data were available until the recovery of the UVC.
- 2014-1-6 11:03:00 ~ 2014-01-06 15:52:00(708036020CENTAURUS_A)
Failure of the star tracking by STT-A (EC 63).
2013
- 2013-12-31 01:30:00 ~ 2014-01-03 15:30:00 (808089030ABELL_1689_-OFFSET)
Failure of the star tracking by STT-A (EC 63).
- 2013-12-25 18:55:00 ~ 2013-12-26 14:30:00 (708022010NGC5929)
Failure of the star tracking by STT-A (EC 63).
2013-12-17 08:39:00~ 2013-12-19 12:04:00 (808053010NGC_253)
Failure of the star tracking by STT-A (EC 63).
- 2013-08-29 21:19:00 ~ 2013-08-31 23:07:00 (508074010MBM36-OFF)
Failure of the star tracking by STT-A (EC 63).
- 2013-08-16 19:25:00 ~ 2013-08-18 12:00:00 (508081010MBM20)
Failure of the star tracking by STT-A (EC 63).
- 2013-08-15 04:02:00 ~ 2013-08-15 10:14:00 (708036010CENTAURUS_A)
Failure of the star tracking by STT-A (EC 63).
- 2013-07-27 10:14 ~ 2013-07-28 12:41 (808104010 A3528-S_OFFSET 1)
Bad tracking of the star by STT-A.
- 2013-07-28 12:41 ~ 2013-07-29 09:33 (808105010 A3528-S_OFFSET 2)
Bad tracking of the star by STT-A.
- 2013-07-30 15:39 ~ 2013-07-31 09:46 (808107010 A3528-S_OFFSET 4)
Bad tracking of the star by STT-A.
- 2013-07-02 7:15 ~ 2013-07-02 16:45 (808090010 COMA W SHOCK1)
Failure of the star tracking by STT-A.
- 2013-04-26
DR did not respond to commands. DR was rebooted
on 2013-04-27 02:10. The data between 2013-04-25 09:00 and
2013-04-27 02:10 was lost.
2012
- 2012-08-11 ~ 2012-08-13 (507037010 G298.6-0.0)
Failure of the star tracking by STT-A.
- 2012-07-31 ~ 2012-08-01 (507076010 MBM16)
Failure of the star tracking by STT-A.
- 2012-07-12
Discharge of RCS propellant are performed. Scientific observation
was halted between 2012-07-12 21:30 and 2012-07-14 03:45.
The orbital elements are measured and updated.
- 2012-06-25 08:20
IRU-S2 OFF.
- 2012-06-25 06:40:
The three-axis attitude control system started referring to the combination of the IRU-X,Y,Z.
- 2012-06-09 14:46
IRU-S1 OFF.
- 2012-06-08 19:45
IRU-S1 ON.
- 2012-01-26
The three axis control by the AOCS was restarted.
The satellite bus system is in a normal operation mode.
- 2012-01-24
BCCU commands the Under Voltage Control (UVC) mode.
The attitude control system was automatically switched into the safehold mode.
The mission instruments XIS and HXD were also shutdown.
No scientific data were available until the recovery of the UVC.
- 2012-01-22 ~ 2010-01-24 (WKK 4438)
Failure of the star tracking by STT-B. The observation was terminated by switching into the UVC mode.
2011
- 2011-11-05 PASS #1 (PKS 0625-354)
The V/T curve parameter for the Battery Current Control Unit (BCCU) was changed from No.4 (46.56V) to No.5(47.06V).
- 2011-07-22 ~ 2011-08-09
IRU-S1 ON. Diagnostic test of its electric outputs.
2010
- 2010-09-29 ~ 2010-10-02 (MAXI J1659-152)
Failure of the star tracking by STT-A.
It is because three of five candidates are too dim to be stably tracked.
HK data show that STT was hardly operated with the star tracking mode
during all the three observations of MAXI J1659-152.
The users are recommended to apply the attitude correction tool
'aeattcorr' for the XIS data to remake the attitude file
during the following durations:
MAXI J1659-152 2010-09-29 05:40:00 -- 2010-09-29 16:37:00
MAXI J1659-152 2010-09-30 05:25:00 -- 2010-09-30 23:30:00
MAXI J1659-152 2010-10-01 08:50:00 -- 2010-10-02 05:32:00
-2010-10-02 (MAXI J1659-152, GX1+4)
A part of DR data was not down-linked and lost.
-2010-07-13 (NGC 4945)
A part of DR data was not down-linked and lost.
-2010-06-19 (E0102-72)
A thuderbold struck the 20m antenna of USC tracking station.
20m, and number of contact pass of Suzaku was reduced from the schedule.
A part of DR data was not down-linked. (E0102-72_PSUM, and Abell 2537)
Some of usual calibration (PSD OFF/ON etc) were skipped.
-2010-06-15 (BLOWOUT_SHELL_2)
IRU-Y ON, since then, the attitude determination on ground can use
IRU X, Y, and Z. (for details, please see JX-ISAS-SUZAKU-MEMO-2010-04)
-2010-05-23 (Mrk 205)
Due to the trouble of the X-band receiver of USC tracking station,
a part of DR data was not down-linked. (Mrk 205 and Fairall 9)
-2010-04-08 (GALACTIC_BULGE5)
Due to the mis-opearation triggered by the S-band receiver
of USC tracking station trouble, DHU was stopped and OP did
not run between 2010/04/08 00:18 and 2010/04/09 00:16.
2009
-2009-12-18 (FESI1)
The combination of IRUs for the attitude control was changed from X-Z-S1 to X-Z-S2 at 03:21 on 18th, December 2009.
-2009-12-14 (SEP #3)
An "abnormal angular momentum change rate" error was detected on
December 14th ~14:00. Calibration by the STT was deactivated and was not
reactivated at December 15th ~01:45.
-2009-12-07 (NEP #2)
An "abnormal angular momentum change rate" error was detected on
December 7th ~21:00. Calibration by the STT was deactivated and was not
reactivated at December 8th ~07:05. The ATOG flag was also detected at December 7th ~21:07:37.
-2009-11-29 ~ 2009-12-04
(A1246_OFF, VELA_SOUTHWEST_1, BLOWOUT6, CYGNUS_X-1, E0102-72_PSUM, MCG-2-58-22)
The HK data was not recorded properly from 2009-11-29 10:00 to 2009-12-04 10:32.
The HK data were referred in the attitude correction tool "aeattcor".
This tool calculates the effects of the "thermal wobbling" caused by thermal distortions of the satellite bodies, and correct the Euler angles in attitude file for Suzaku. Unfortunately, the attitude during the days were not processed with the tool "aeattcor" properly.
-2009-11-28 (A1246 OFFSET)
An "abnormal angular momentum change rate" error was detected on
November 28th ~15:50. Calibration by the STT was deactivated and was not
reactivated at November 28rd ~17:30.
An "abnormal angular momentum change rate" error was again detected on
November 28th ~22:01. Calibration by the STT was deactivated and was not
reactivated until the end of the observation at November 29th ~23:03.
-2009-11-19 (RIM3)
An "abnormal angular momentum change rate" error was detected on
November 11th ~15:59. Calibration by the STT was deactivated and was not
reactivated at November 19th ~16:10.
-2009-11-03 (NGC_4138)
An "abnormal angular momentum change rate" error was detected on
November 3rd ~09:15. Calibration by the STT was deactivated and was not
reactivated until the end of the observation at November 3rd ~21:14.
-2009-10-13 (GALACTIC_BULGE7)
An "abnormal angular momentum change rate" error was detected on
October 13th ~17:38. Calibration by the STT was deactivated and was not
reactivated until the end of the observation at October 14th ~11:29.
-2009-09-16 (GALACTIC_BULGE8)
An "abnormal angular momentum change rate" error was detected on
September 16th ~05:03. Calibration by the STT was deactivated and was
not reactivated until the end of the observation at September 16th ~07:18.
-2009-09-15 (GALACTIC_BULGE8)
An "abnormal angular momentum change rate" error was detected on
September 15th ~18:05. Calibration by the STT was deactivated and was
not reactivated until September 16th ~01:00.
-2009-09-09 (HESSJ1809-193A)
An "abnormal angular momentum change rate" error was detected on
September 9th ~23:11. Calibration by the STT was deactivated and was
not reactivated until September 10th ~03:01.
-2009-09-08 (ASO 0376)
An "abnormal angular momentum change rate" error was detected on
September 8th ~06:45. Calibration by the STT was deactivated and was
not reactivated until the end of the observation at September 9th
~02:40. The maneuver to the next target 4U1820-30 was forced to be
delayed from September 8th ~13:50 to September 9th ~02:40.
-2009-08-29 (KEPLER_BG_GE)
An "abnormal angular momentum change rate" error was detected on
August 29 ~00:16. Calibration by the STT was deactivated and was not
reactivated until August 29 ~08:09.
-2009-08-28 (PERSEUS_1_4_WIN)
An "abnormal angular momentum change rate" error was detected on
August 28 ~03:27. Calibration by the STT was deactivated and was not
reactivated until August 28 ~09:52
-2009-06-27 (A1795_NEAR_WEST)
The HXD-DE was halted during a SAA passage.
The recovery operation was performed on July 1.
There is no telemetry data of HXD between June 27 17:27 - July 1 ~16:30.
-2009-06-23 (AO_PSC)
A trouble was occurred in XIS0 on June 23 02:01.
See the detail page.
-2009-06-16 (CYGNUS_LOOP_P8 / NGC6888) An "abnormal angular momentum
change rate" error was detected on June 16 ~22:27. Calibration by the
STT was deactivated and was not reactivated until June 17 ~18:27. Note
that the maneuver to the next target (NGC6888) was made at June 17
01:36.
- 2009-05-10 (LockmanHole)
SAFEHOLD FROM 2009-05-10 11:13 TO 2009-05-12 14:15
This safehold was caused by an exceed of rotation speed in one of momentum
wheels, during the maneuver of spacecraft.
2008
- 2008-11-25 (NGC 4051, NGC 3556)
Data recorder became full and stopped from 2008-11-25 13:44 to 2008-11-25 23:24.
No science/HK data was recorded from 2008-11-25 13:44 to 2008-11-25 23:24.
2007
- 2007-12-08
Trouble in the on-board data processing of XIS0.
No usable data were output for Segments C & D of XIS0 due to the bit-upset
in one of the data processing units of the sensor.
No data for XIS0, Seg. C & D from 2007-12-8 4:48 to 12-9 1:16 (NGC4636 North)
No data for XIS0, Seg. C & D from 2007-12-9 1:16 to 12-11 1:10 (Swift J2127.4+6564)
No data for XIS0, Seg. C & D from 2007-12-11 1:10 to 12-11 20:30 (3EG 1234+1318 3)
No data for XIS0, Seg. C & D from 2007-12-11 20:30 to 12-12 0:30 (3EG 1234+1318 4)
- 2007-09-15
Operations for Pass 1-5 were canceled due to the typhoon attack.
Data recorder became full at around the end of following
three days, and stopped from 2007-09-17 16:47 to 2007-09-17 17:30 (MCG+8-11-11).
- 2007-08-02
Operations for Pass 1-5 were canceled due to the typhoon attack.
No affection to the observation data.
- 2007-07-14
Operations for Pass 1-5 were canceled due to the typhoon attack.
Data recorder became full at around the end of following three days,
and stopped from 2007-07-15 21:56 to 2007-07-16 01:35 (NGC5548 and NGC1052).
- 2007-07-01 (3C273, NGC5548, TW_HYA)
A generic protection error is occurred during the SAA passage,
and the DP switched into the hardware backup mode.
No observation data is recorded from 2007-07-01 06:23 to 2007-07-02 13:45.
- 2007-06-05
STT reference star catalogs are updated. (on ground)
329 stars were removed. 25 are too bright while 304 are too faint.
426 variable stars (Class 2) are newly added.
- 2007-06-01 (NGC 4395)
AOCS ERROR CODE 63
Failure of the star tracking by STT-A. It is because two of four candidates
are too dim to be tracked sometimes.
- 2007-04-18
The solid neon of the XRS was gone on April 18, 2007 (UT), and temperatures
around the neon tank has begun to gradually increase. The neon lifetime is
consistent with the estimation after the helium loss, without helium vapor
cooling.
- 2007-03-05 (GC_SgrB_East)
AOCU ERROR CODE 63 was detected, which means STT was not used to control
attitude for more than 3 revolutions. This was caused by an inappropriate
guide star selections of STT. HK data show that STT was used for at least
some time during this observation, and hence impact on the observations is
considered small. The problematic two stars were eliminated from the
starcatalog hereafter.
- 2007-12-20 (XBONGBLUE1)
The DHU was running with the AOCU dump mode during 00:14:40-01:53:10,
in which AOCU data was not contained.
SAFEHOLD FROM 2007-02-20 04:11 TO 2007-02-21 02:34
Planned safehold to update the AOCU onboard software in EPROM. Blank sky
was observed from 2007-02-21 02:34 to 2007-02-21 04:11.
2006
- YEAR-END AND NEW-YEAR OPERATION (NGC 4649)
High-voltage of HXD was set to zero at 2006-12-30 05:04. Returned to
normal value at 2007-01-04 03:16.
- 2006-11-18 (Mrk 766)
SAFEHOLD FROM 2006-11-18 08:43 TO 2006-11-21 00:33
This safehold was caused by a wrong parameter setting for the attitude
control subsystem.
- 2006-11-09 (Draco Enhancement)
XIS-2 anomaly occurred.
- 2006-09-08 Pass #1 (Galactic Center)
IRU-SB Loop on. AOCU ERROR CODE 63 was detected, which actually
occurred during the observation of T CRB. See description on
2006-02-22. Further investigation is underway.
- 2006-09-07 Pass #2 (T CRB)
Heater of IRU-SB was turned on.
- 2006-09-07 Pass #1 (T CRB)
Pass #1 was cancelled due to conflict with Solar-B ground test. As
a result, DR was operated with low rate during the following period.
DATA RATE LOW FROM 2006-09-07 8:42(TBR) TO 10:12(TBR) (T CRB)
- 2006-09-06 Pass #1 (Crab offset)
Pass #1 was cancelled due to conflict with Solar-B ground test. As
a result, we could not upload the attitude commands for Crab offset P10,
and Suzaku pointed to E0102.2-7219. DR was operated with low
rate between pass 1 and 2.
WRONG POINTING FROM 2006-09-06 11:40 TO 13:20 (CRAB OFFSET P10)
DATA RATE LOW FROM 2006-09-06 10:25(TBR) TO 11:55(TBR) (CRAB OFFSET P10)
- 2006-09-05 Pass #2, #4, #5 (Crab)
Pass #2 was cancelled due to thunder, and pass #4 and #5 were
cancelled due to power failure. These problems prevented us to upload some of the
attitude commands for the Crab offset pointings (P6, 7, 8, 9). Because of this,
Suzaku was pointed at NGC1667, SNR0509-67.5, A399_401_filament,
Sigma2_CRB.
WRONG POINTING FROM 2006-09-06 05:20 TO 07:00 (CRAB OFFSET P6)
WRONG POINTING FROM 2006-09-06 07:00 TO 08:30 (CRAB OFFSET P7)
WRONG POINTING FROM 2006-09-06 08:30 TO 10:10 (CRAB OFFSET P8)
WRONG POINTING FROM 2006-09-06 10:10 TO 11:40 (CRAB OFFSET P9)
- 2006-07-05 Pass #1, #2, #3, #5 (NGC5044 offset 2)
Pass #1, #2, #3, and #5 were cancelled due to thunder. Only Pass #4
was performed. As a result, DR was operated with low rate during the
following period:
DATA RATE LOW FROM 2006-07-05 16:29 TO 2006-07-05 21:19
- TARGET OUT OF FOV FROM 2006-08-01 03:15 TO 2006-08-01 09:58 (1E1207.4-5209)
- TARGET OUT OF FOV FROM 2006-08-01 09:58 TO 2006-08-02 05:35 (Mrk 1073)
These were caused by the particle events in STT during SAA. A trimming
maneuver was performd at 2006-08-02 05:35.
- TARGET OUT OF FOV FROM 2006-08-09 23:41 TO 2006-08-10 00:07 (NGC5506)
- SAFEHOLD FROM 2006-08-10 00:07 TO 2006-08-11 02:07
These were caused by the particle events in STT during SAA. Now STT is
explicitly disabled during SAA to prevent similar problems.
- 2006-06-15 Pass #1, #2 (NGC3923)
Pass #1 and #2 were cancelled due to strong wind. As a result, DR was
operated with low rate during the following period:
DATA RATE LOW FROM 2006-06-15 01:20 TO 2006-06-15 04:29
OP START at Pass #3 was delayed by about 40 sec.
- 2006-05-24 Pass #1 (1ES 1959)
IRU-SB loop and heater were turned off.
- 2006-05-16 (Beta Lyra 3)
OP was uploaded at Pass #2, but Pass #1-#2 was not taken care of by
the previous OP. As a result, DR was operated with low rate during the
following period:
DATA RATE LOW FROM 2006-05-16 15:26 TO 2006-05-16 16:53
- 2006-05-07/08 PASS #1-#5 (Beta Lyra 1, M81)
Problem of the time assignment equipment of the 34m antenna, due to bad
connection of one of the connectors.
- 2006-04-30/2006-05-01 (1RXS J213344.1+510)
The limit to the DR operation was further relaxed and now we use data
rate high during the contact orbits. It is almost the same as that
before the DR problem occurred, except for the pointer controls.
- 2006-04-25 PASS #1 (LMC X-2)
Cancelled due to hardware failure of the antenna drive. OP was
uploaded at pass #2. As a result, DR was operated with low rate during
the following period:
DATA RATE LOW FROM 2006-04-25 01:59 TO 2006-04-25 03:27
- 2006-04-22 (LMC_X-3_OFF)
The limit to the DR operation was slightly relaxed and now we use data
rate medium during the earth occultation if possible.
- 2006-04-20 PASS #1, #3, #5 (1E1841-045)
Cancelled due to strong wind. OP was uploaded at pass #2. As a result,
DR was operated with low rate during the following period:
DATA RATE LOW FROM 2006-04-20 05:28 TO 2006-04-20 06:55
- 2006-04-14 PASS #5 (IRAS08572+3915)
Recording and reproducing pointers of the DR were set to 00000 and A0000 after
all the data were played back, in order for the recording pointer not to step
over the border BFFFF-00000, and not to catch up with the reproducing pointer.
- 2006-04-14 (IRAS08572+3915)
To minimize the data loss due to the DR problem, we started to reduce
the DR usage as much as possible. Now the data rate low is used during
the earth occultataion.
- 2006-04-06 PASS #5 (HESS_J1804)
Playback error of DR, when the reproducing pointer returned from BFFFF
to 00000. After this event, DR sometimes caused a problem until April
14, when we introduced pointer controls.
- DATA LOST FROM 2006-04-10 04:55 TO 2006-04-10 08:47 (G11.2-0.3)
- DATA LOST FROM 2006-04-10 08:57 TO 2006-04-10 10:27 (G11.2-0.3)
These were caused by irregular stop of DR recording.
- DATA LOST FROM 2006-04-12 21:56 TO 2006-04-13 01:25 (SWIFT J2000.6+3210)
- DATA LOST FROM 2006-04-13 01:25 TO 2006-04-13 12:06 (SWIFT J0350.1-5019)
- DATA LOST FROM 2006-04-13 12:18 TO 2006-04-13 13:48 (SWIFT J0350.1-5019)
These were caused by irregular stop of DR recording.
- SAFEHOLD FROM 2006-03-25 10:38 TO 2006-03-27 22:40
This safehold was caused by the particle event of the solar sensor.
Operational countermeasure was adopted.
- PROBLEM OF TIME ASSIGNMENT IN 20M SYSTEM FROM 2006-03-21 TO 2006-03-31
The time assignment equipment for the 20m system was refurbished on
March 21. However, it was not perfect. This problem was fixed on March
31, but there was a time assignment error for the real data during
this period.
- 2006-03-19/20 PASS #1 (VICINITY OF LMC X-3)
AOCU ERROR CODE 63 was detected. See description on 2006-02-22. Star
trackers were available only for a small portion of time, and hence,
the absolute pointing accuracy was slightly worse.
- 2006-02-22 PASS #1 and PASS #4 (SGR_C)
AOCU ERROR CODE 63 was detected, which means STT was not used to contro
attitude for more than 3 revolutions. This was caused by an inappropriate
guide star selections of STT. HK data show that STT was used for at least
some time during this observation, and hence impact on the observations is
considered small.
- 2006-02-20, 2006-02-21 (HIGH_LATTITUDE_B, SGR_C)
No Suzaku operations due to the launch of ASTRO-F (Akari).
2005
- YEAR-END AND NEW-YEAR OPERATION (NGC720)
High-voltage of HXD was set to zero at 2005-12-30 12:19. Returned to
normal value at 2006-01-04 12:45.
- SAFEHOLD FROM 2005-12-14 12:23 TO 2005-12-15 23:40
This safehold was triggered due to the large deviation of the satellite's angular
momentum from the nominal direction, which was caused by a large gravitational
gradient during the maneuver. From now on, accumulation of the gravitational torque
will be checked before maneuver commands are uploaded.
- 2005-11-11 PASS #3 (N4051)
Data were lost for about 30 sec at 13:42, due to the poor contact to the
satellite.
- 2005-11-07 (NGC2992, A3376)
No data were played back due to the trouble in the ground station.
At pass #2, new OP was uploaded to reduece the DR usage (weekend-type
telemetry distribution from 2005-11-07 13:57 to 2005-11-08 13:57).
However, DR became full at around 2005-11-07 23:34. As a result, data
during the following period were completely lost:
DATA LOST FROM 2005-11-07 23:34 TO 2005-11-08 12:07 (A3376)
- 2005-10-30/31 PASS #2 AND #3 (GALACTIC_RIDGE)
Discontinuity was detected in the playbacked data, which means a small
amount of data were lost when recording.
- SAFEHOLD FROM 2005-10-19 22:35 TO 2005-10-21 01:53
Planned safehold to update the AOCU onboard software in RAM. Blank sky
and maneuver tests from 2005-10-21 01:53 to 2005-10-22 01:55.
- 2005-10-12 PASS #1 (GC_SRC_B2)
Problem of command transmitter, and OP could not be uploaded. As a
result, DR was operated with data rate L during the following period:
DATA RATE LOW FROM 2005-10-12 2:11 TO 2005-10-12 3:41
- 2005-09-08 (SN1006_NE_BG)
Small amount of the data were lost due to the problem of the demodulator
in the ground station.
- 2005-09-05, 2005-09-06, 2005-09-07 (SN1006_NE_BG)
On Sep 5 and 6, all the contact passes (#1-#5) were cancelled due to a
big typhoon. On Sep 7, passes #1-#4 were cancelled due to electric
power failure. At pass #5, DR (data recorder) was played back, but we
could not upload OP (operation program). As a result, DR was operated
with low rate during the following period:
DATA RATE LOW FROM 2005-09-04 15:00 TO 2005-09-07 02:13
DATA RATE LOW FROM 2005-09-07 02:18 TO 2005-09-07 17:43
- 2005-08-23 PASS #1 (NGC4945)
Onboard alignment matrix was updated so that the spacecraft Z axis was
aligned with the XRT FOV. This affects the spacecraft Euler angles,
and the onboard attitude determination, but does not affect
the ground attitude determination results.